# Koby Hayashi

### v1

@@ 3.1 SymNMF with Low-rank Approximate Input
The computational complexity of Algorithm LAI-SymNMF is that of the RRF followed by cost for iteratively updating the factors. Again, the cost of the RRF is $O(qm^{2}l)$. Then computing ${\bm{\mathbf{{V}}}}={\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{U}}}}$, where ${\bm{\mathbf{{U}}}}$ is the output of the RRF, ${\bm{\mathbf{{V}}}}$ and ${\bm{\mathbf{{U}}}}$ are $m\times l$ and ${\bm{\mathbf{{X}}}}={\bm{\mathbf{{X}}}}^{\sf T}$ is $m\times m$, costs $O(m^{2}l)$. Additionally, each iteration requires forming two Gram matrices costing $O(mk^{2})$ each and applying the LAI to the factor matrices costing $O(mkl)$. If the algorithm runs for $t$ iterations then the overall cost is $O(qm^{2}l+tmkl)$ and $l\geq k$. So if $tk\ll qm$ then we expect that computing the low-rank approximate input via the RRF will dominate the run time. Naturally, the choice of update function will determine the update cost.
@@ 3.2 Approximation Errors for LAI-NMF
The authors of [48] presented a simple error bound applicable to LAI-NMF that we adapt to our needs. Proposition 1 from [48] states the following:
Proposition 1.
@@
$\upsilon^{*}\leq\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}^{*}({\bm{\mathbf{{H}}}}^{*})^{\sf T}\|_{F}\leq 2\mu+\upsilon^{*}$ (22)
Proof.
@@
Proposition 1 allows us to reason about the achievable quality of approximation of LAI-NMF. Choosing a larger $l$ can help decrease $\mu$ but will also result in higher computational complexity. A natural choice for computing ${\bm{\mathbf{{U}}}}_{X}$ and ${\bm{\mathbf{{V}}}}_{X}$ is the truncated SVD (which would minimize $\mu$) or, as we use, an approximate truncated SVD or EVD computed using the RRF. As an alternative, the intermediate inequality, from the proof of Proposition 1,
$\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}^{*}({\bm{\mathbf{{H}}}}^{*})^{\sf T}\|_{F}\leq\mu+\|{\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}-{\bm{\mathbf{{W}}}}_{+}({\bm{\mathbf{{H}}}}_{+})^{\sf T}\|_{F}$ (23)
@@
Proposition 1 can give an error bound for Algorithm LAI-SymNMF. For this, one needs an error bound for the QB-decomposition or the approximate EVD. That is, given a decomposition from the RRF as ${\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}$, we seek a value $\mu_{RRF}\geq\|{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}-{\bm{\mathbf{{X}}}}\|_{F}$. Theorem 11, whose statement can be found in Appendix H, from [17] provides such a bound. That is, $\mu_{RRF}\geq\|{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}-{\bm{\mathbf{{X}}}}\|_{F}$ holds with some chosen probability $1-\delta$, where $\mu_{RRF}$ depends on $\delta$ and other parameters of the RRF such as $l$ and $q$. A full statement of this can be found in Section A.1. Given $\mu_{RRF}$, one can obtain an error bound for the approximate, truncated EVD produced by Algorithm Apx-EVD. We use a fact from [18]. Given a low-rank approximation ${\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}\approx{\bm{\mathbf{{X}}}}$, where ${\bm{\mathbf{{X}}}}$ is symmetric, from Algorithm RRF and defining ${\bm{\mathbf{{P}}}}_{X}={\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{Q}}}}_{X}^{\sf T}$, observe that
@@ 3.3 Practical Considerations for LAI-SymNMF
Iterative Refinement (IR) runs some number of NMF iterations using the full matrix ${\bm{\mathbf{{X}}}}$ instead of the LAI. That is, after the iterations of LAI-SymNMF are finished, the algorithm switches over to using the full ${\bm{\mathbf{{X}}}}$ matrix, therefore capturing information possibly lost in the low-rank approximation of ${\bm{\mathbf{{X}}}}$. This helps in cases where the right side of Equation 23 is large. In practice our experimental results show that this method is effective in improving the SymNMF approximations attained by LAI-SymNMF while running faster than standard SymNMF methods.
@@
The RRF has two main hyperpameters 1) column oversampling parameter $\rho$ and 2) the power iteration parameter $q$. There exists work on adaptive methods for selecting $\rho$ [18]. For our algorithms, where $k$ is usually considered a static input to NMF methods, we find that choosing $\rho$ in the range of $2k$ to $3k$ is satisfactory. Empirically we find that determining a good $q$ is more difficult. Prior works recommend a choice of $q=2$ [13, 18]. However we find that this choice can be inadequate and negatively impact performance. To remedy this we propose an Adaptive RRF algorithm that automatically chooses $q$. This method checks the residual of the QB-Decomposition after each power iteration and stops once a certain stopping criteria is met (e.g. lack of reduction in residual), similar to NMF. The residual check is cheap. Checking the residual of the QB-Decomposition after each power iteration requires only one extra multiplication against ${\bm{\mathbf{{X}}}}$ when calling the RRF by use of a standard ‘trick’ for computing the residual. That is if $q$ power iterations are performed we only apply ${\bm{\mathbf{{X}}}}$, $q+1$ times. Pseudocode can be found in the Appendix in Algorithm Ada-RRF. This approach ensures we achieve a good value of $\mu_{RRF}$ as in Equation 23.
@@ 3.4 Discussion of LAI-SymNMF
Compared to existing randomized methods, such as those in [13, 41], the LAI method is more general in that it can work for any NMF method that relies on matrix vector products ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{v}}}}$, where ${\bm{\mathbf{{v}}}}$ is an arbitrary vector, for performance. For example, the Compressed-NMF method from [41], which we compare against in Section 5, is only applicable for Alternating Updating methods. The PGNCG method for SymNMF from [15] can be used for LAI-SymNMF but not for Compressed-NMF. Finally, unlike the randomized method in [13], the LAI method decouples the randomization from the NMF algorithm and accordingly preserves the convergence properties, such as convergence to a stationary point [23], of existing NMF algorithms applied to the low-rank input and can be reasoned about via Proposition 6. A more detailed discussion of existing randomized NMF methods and LAI-NMF is given in Section A.2.
@@ 4.1 Leverage Score Sampling for Multiple Right hand Sides
Unlike in the CP decomposition, the coefficient matrix in the LS problem for low-rank matrix approximation has, in general, no special structure we can exploit to obtain fast leverage score estimates. However, for many methods the products ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ and ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ are the most expensive part of an NMF iteration[21]. By computing a thin ${\bm{\mathbf{{Q}}}}{\bm{\mathbf{{R}}}}$ factorization of the matrices ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ at each iteration we can obtain exact leverage scores for use in sampling and avoid the expensive full matrix products involving the data matrix ${\bm{\mathbf{{X}}}}$. Pseudocode for the algorithm is given in Algorithm LvS-SymNMF.
@@
$\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}-{\bm{\mathbf{{X}}}}\|_{F},$ (24)
where the coefficient matrix ${\bm{\mathbf{{W}}}}\in\mathbb{R}^{m\times k}$ is much smaller than the right hand side matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times m}$ if $k\ll n$. Consider (approximately) solving the problem by the update as ${\bm{\mathbf{{H}}}}:=\text{Update}({\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}},{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{W}}}})$. Recall Update() was introduced for Algorithm LAI-SymNMF and its details can be found in Appendix G. Computing ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ costs $O(mk^{2})$ and $O(m^{2}k)$ flops. The cost of the Update() will be denoted as $O(T(m,k))$ and is dependent on the method used. In light of this we suggest the following randomized approach:
@@
    Perform an Update() for the reduced problem $\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}-{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{X}}}}\|_{F}^{2}$ as ${\bm{\mathbf{{H}}}}:=\text{Update}({\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{S}}}}_{W}^{\sf T}{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{W}}}},{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{S}}}}_{W}^{\sf T}{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{W}}}})$.
    (As opposed to ${\bm{\mathbf{{H}}}}:=\text{Update}({\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}},{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{W}}}})$.)
The conditions for this scheme to provide speed up are roughly that $s\ll m$ and that the cost $T(m,k)$ does not dominate the overall complexity. The key observation here is that computing the thin QR-decomposition costs only $O(mk^{2})$ flops and so when a large right hand side (RHS) is present, computing the leverage scores is not the dominating cost. Note that this observation is relevant for problems with a similar structure to Equation 24. For example one can approximately solve an OLS problem with a many RHS vectors and small coefficient matrix using this scheme.

### v2

@@ 3.1 SymNMF with Low-rank Approximate Input
The major part of computational complexity of Algorithm LAI-SymNMF is due to the RRF and iteratively updating the factors. Again, the cost of the RRF is $O(qm^{2}l)$. Then computing ${\bm{\mathbf{{V}}}}={\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{U}}}}$, where ${\bm{\mathbf{{U}}}}$ is the output of the RRF, ${\bm{\mathbf{{V}}}}$ and ${\bm{\mathbf{{U}}}}$ are $m\times l$ and ${\bm{\mathbf{{X}}}}={\bm{\mathbf{{X}}}}^{\sf T}$ is $m\times m$, costs $O(m^{2}l)$. Additionally, each iteration requires forming two Gram matrices costing $O(mk^{2})$ each and applying the LAI to the factor matrices costing $O(mkl)$. If the algorithm runs for $t$ iterations then the overall cost is $O(qm^{2}l+tmkl)$ and $l\geq k$. So if $tk\ll qm$ then we expect that computing the low-rank approximate input via the RRF will dominate the run time. Naturally, the choice of update function will determine the update cost.
@@ 3.2 Approximation Errors for LAI-NMF
The authors of [59] presented a simple error bound applicable to LAI-NMF that we can use to reason about LAI-NMF’s performance. Proposition 1 from [59] states the following:
Proposition 3.1.
@@
(16) $\upsilon^{*}\leq\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}^{*}({\bm{\mathbf{{H}}}}^{*})^{\sf T}\|_{F}\leq 2\mu+\upsilon^{*}$
Proof 3.2.
@@
Proposition 3.1 allows us to reason about the achievable quality of approximation of LAI-NMF. Choosing a larger $l$ can help decrease $\mu$ but will also result in higher computational complexity. A natural choice for computing ${\bm{\mathbf{{U}}}}_{X}$ and ${\bm{\mathbf{{V}}}}_{X}$ is the truncated SVD (which would minimize $\mu$) or, as we use, an approximate truncated SVD or EVD computed using the RRF. As an alternative, the intermediate inequality, from the proof of Proposition 3.1,
(17) $\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}^{*}({\bm{\mathbf{{H}}}}^{*})^{\sf T}\|_{F}\leq\mu+\|{\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}-{\bm{\mathbf{{W}}}}_{+}({\bm{\mathbf{{H}}}}_{+})^{\sf T}\|_{F}$
@@
Proposition 3.1 can give an error bound for Algorithm LAI-SymNMF. For this, one needs an error bound for the QB-decomposition or the approximate EVD. That is, given a decomposition from the RRF as ${\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}$, we seek a value $\|{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}-{\bm{\mathbf{{X}}}}\|_{F}\leq\mu_{RRF}$. Theorem 3.3, from Gu [25], provides such a bound. Theorem 3.3 is a partial statement of Theorem 5.8 from [25] that we include for reference and completeness.
Theorem 3.3.
Let ${\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}$ be a low-rank approximation of ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times n}$, with $n\leq m$, obtained from the RRF with desired low-rank $r$, power iteration parameter $q$, column over sampling parameter $\rho=l-r$ and a parameter $10<\delta<<$ 1. Define
$C_{\delta}=\frac{e\sqrt{l}}{\rho+1}\left(\frac{2}{\delta}\right)^{\frac{1}{\rho+1}}\Big{(}\sqrt{n-l+\rho}+\sqrt{l}+\sqrt{2\log(2/\delta)}\Big{)}.$
Then with probability $1-\delta$ the following holds:
$\|{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}-{\bm{\mathbf{{X}}}}\|_{F}\leq\sqrt{\Big{(}\sum_{j=r+1}^{n}\sigma_{j}^{2}({\bm{\mathbf{{X}}}})\Big{)}+rC_{\delta}^{2}\sigma^{2}_{r+1}({\bm{\mathbf{{X}}}})\Big{(}\frac{\sigma_{r+1}({\bm{\mathbf{{X}}}})}{\sigma_{r}({\bm{\mathbf{{X}}}})}\Big{)}^{4q}}.$
This Theorem says that $\|{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}-{\bm{\mathbf{{X}}}}\|_{F}\leq\mu_{RRF}$ holds with some chosen probability $1-\delta$, where $\mu_{RRF}$ depends on $\delta$ and other parameters of the RRF such as $l$ and $q$.
Proposition 3.4 explicitly combines Proposition 3.1 and Theorem 3.3 to give a probabilistic error bound for LAI-NMF with a randomized low-rank input from the RRF.
Proposition 3.4.
Given a matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times n}$ compute a low-rank approximation ${\bm{\mathbf{{X}}}}\approx{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}$ where ${\bm{\mathbf{{Q}}}}_{X}\in\mathbb{R}^{m\times l}$ and ${\bm{\mathbf{{B}}}}_{X}\in\mathbb{R}^{l\times n}$ from the RRF. Then by Theorem 3.3, for any $0<\delta\ll 1$, we have that
$\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}\|_{F}\leq\sqrt{\Big{(}\sum_{j=k+1}^{n}\sigma_{j}^{2}({\bm{\mathbf{{X}}}})\Big{)}+kC_{\delta}^{2}\sigma^{2}_{k+1}({\bm{\mathbf{{X}}}})\Big{(}\frac{\sigma_{k+1}({\bm{\mathbf{{X}}}})}{\sigma_{k}({\bm{\mathbf{{X}}}})}\Big{)}^{4q}}=\mu_{RRF}$
holds with probability $1-\delta$. Define $\{{\bm{\mathbf{{W}}}}^{*},{\bm{\mathbf{{H}}}}^{*}\}=\operatorname*{arg\,min}_{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}$ and the optimal NMF error $\upsilon^{*}=\min_{\{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\}\geq 0}\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F}$. Then with probability $1-\delta$
(18) $\upsilon^{*}\leq\|{\bm{\mathbf{{X}}}}-{\bm{\mathbf{{W}}}}^{*}({\bm{\mathbf{{H}}}}^{*})^{\sf T}\|_{F}\leq 2\mu_{RRF}+\upsilon^{*}$
as in Proposition 3.1.
In the case that ${\bm{\mathbf{{X}}}}$ is symmetric and the RRF is used to compute an approximate EVD it is simple to extend Proposition 3.4. Given $\mu_{RRF}$, one can obtain an error bound for the approximate, truncated EVD produced by Algorithm Apx-EVD. We use a fact from [26]. Given a low-rank approximation ${\bm{\mathbf{{X}}}}\approx{\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}$, where ${\bm{\mathbf{{X}}}}$ is symmetric, from Algorithm RRF and defining ${\bm{\mathbf{{P}}}}_{X}={\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{Q}}}}_{X}^{\sf T}$, observe that
@@ 3.3 Practical Considerations for LAI-SymNMF
Iterative Refinement (IR) runs some number of NMF iterations using the full matrix ${\bm{\mathbf{{X}}}}$ instead of the LAI. That is, after the iterations of LAI-SymNMF are finished, the algorithm switches over to using the full input matrix ${\bm{\mathbf{{X}}}}$, therefore capturing information possibly lost in the low-rank approximation of ${\bm{\mathbf{{X}}}}$. This helps in cases where the right side of Equation 17 is large. In practice our experimental results show that this method is effective in improving the SymNMF approximations attained by LAI-SymNMF while running faster than standard SymNMF methods.
@@
The RRF has two main hyperpameters 1) column oversampling parameter $\rho$ and 2) the power iteration parameter $q$. There exists work on adaptive methods for selecting $\rho$ [26]. For our algorithms, where $k$ is usually considered a static input to NMF methods, we find that choosing $\rho$ in the range of $2k$ to $3k$ is satisfactory. Empirically we find that determining a good $q$ is more difficult. Prior works recommend a choice of $q=2$ [19, 26]. However we find that this choice can be inadequate and negatively impact performance. To remedy this we propose an Adaptive RRF algorithm that automatically chooses $q$. This method checks the residual of the QB-Decomposition after each power iteration and stops once a certain stopping criteria is met (e.g. lack of reduction in residual), similar to NMF. The residual check is cheap. Checking the residual of the QB-Decomposition after each power iteration requires only one extra multiplication against ${\bm{\mathbf{{X}}}}$ when calling the RRF by use of a standard ‘trick’ for computing the residual. That is if $q$ power iterations are performed we only apply ${\bm{\mathbf{{X}}}}$, $q+1$ times. The algorithm outline can be found in the Appendix in Algorithm Ada-RRF. This approach ensures we achieve a good value of $\mu_{RRF}$ as in Equation 17.
@@ 3.4 Discussion of LAI-SymNMF
Compared to existing randomized methods, such as those in [19, 51], the LAI method is more general in that it can work for any NMF method that relies on matrix vector products ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{v}}}}$, where ${\bm{\mathbf{{v}}}}$ is an arbitrary vector, for performance. For example, the Compressed-NMF method from [51], which we compare against in Section 5, is only applicable for Alternating Updating methods. The PGNCG method for SymNMF from [22] can be used for LAI-SymNMF but not for Compressed-NMF. Finally, unlike the randomized method in [19], the LAI method decouples the randomization from the NMF algorithm and accordingly preserves the convergence properties, such as convergence to a stationary point [32], of existing NMF algorithms applied to the low-rank input and can be reasoned about via Proposition 3.4. A more detailed discussion of existing randomized NMF methods and LAI-NMF is given in Section B.1.
@@ 4.1 Leverage Score Sampling for Multiple Right Hand Sides
Unlike in the CP decomposition, the coefficient matrix in the LS problem for low-rank matrix approximation has, in general, no special structure we can exploit to obtain fast leverage score estimates. However, for many methods the products ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ and ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ are the most expensive part of an NMF iteration [29]. By computing a thin ${\bm{\mathbf{{Q}}}}{\bm{\mathbf{{R}}}}$ factorization of the matrices ${\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{H}}}}$ at each iteration we can obtain exact leverage scores for use in sampling and avoid the expensive full matrix products involving the data matrix ${\bm{\mathbf{{X}}}}$. The algorithm outline is given in Algorithm LvS-SymNMF.
@@
(19) $\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}-{\bm{\mathbf{{X}}}}\|_{F},$
where the coefficient matrix ${\bm{\mathbf{{W}}}}\in\mathbb{R}^{m\times k}$ is much smaller than the right hand side matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times m}$ if $k\ll n$. Consider (approximately) solving the problem by the update as ${\bm{\mathbf{{H}}}}:=\text{Update}({\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}},{\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}})$. Recall Update() was introduced for Algorithm LAI-SymNMF and its details can be found in Appendix E. Computing ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$ and ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ costs $O(mk^{2})$ and $O(m^{2}k)$ flops. The cost of the Update() will be denoted as $O(T(m,k))$ and is dependent on the method used. In light of this we suggest the following randomized approach:
@@
    Perform an Update() for the reduced problem $\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}-{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{X}}}}\|_{F}^{2}$ as ${\bm{\mathbf{{H}}}}:=\text{Update}({\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{S}}}}_{W}^{\sf T}{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{W}}}},{\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{S}}}}_{W}^{\sf T}{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{X}}}})$.
    (As opposed to ${\bm{\mathbf{{H}}}}:=\text{Update}({\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}},{\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}})$.)
The conditions for this scheme to provide speed up are roughly that $s\ll m$ and that the cost $T(m,k)$ does not dominate the overall complexity. The key observation here is that computing the thin QR-decomposition costs only $O(mk^{2})$ flops and so when a large number of right hand side (RHS) vectors is present, computing the leverage scores is not the dominating cost. Note that this observation is relevant for problems with a similar structure to Equation 19. For example one can approximately solve an OLS problem with many RHS vectors and small coefficient matrix using this scheme.